Given P(A) = 0.40, P(B) = 0.50, P(A ∩ B) = 0.15. Which of the following
is true?
A. A and B are independent
B. A and B are mutually exclusive
C. A and B are complements to each other
D. A and B are not independent
Answer #1

Answer
Option D

A. A and B are independent
to prove it P(A∩ B)=P(A)*P(B) =0.4*0.5=0.2, but its 0.15, so the two are not independent
B. A and B are mutually exclusive
to prove it, P(A∩ B)=0, but its 0.15
C. A and B are complements to each other
To prove it, P(A)+P(B)=1, but it is 0.4+0.5=0.9